Dear Professor,
i would very much appreciate it if you could help me solve the following algebra problem.
thanks
solve the system of equations,using substitiution or eliminationor back-substitution etc.
x+3y-z=13
2x -5z=23
4x-y-2z=4

Answer by richwmiller(17219)

x = -12/17, y = 50/17, z = -83/17
x+3y-z=13
2x -5z=23
4x-y-2z=4
multiply the first by -2
-2x-6y+2z=-26
4x-y-2z=4
add
2x-7y=-22
multiply #1 by 5
5x+15y-5z=65
2x -5z=23
subtract
3x+15y=42
divide by 3
x+5y=14
2x-7y=-22
#1 by 2
2x+10y=28
2x-7y=-22
subtract
17y=50
y=50/17
2x-7y=-22
2x-7*50/17=-22
x-7*25/17=-11
x=-12/17
2x -5z=23
2*(-12/17)-5z=23
-24/17-5z=23*17/17
z=-83/17